A 62.7 kg skier rides a 916.6 m long lift to the top of a mountain. The lift makes an angle of 28.3o with the horizontal. How fast will she be traveling when she reaches the bottom of the mountain?
1
Expert's answer
2020-02-10T09:23:40-0500

If we neglect friction, all we have to do is to use the conservation of energy: potential energy converts to kinetic energy. At the highest point the potential energy is


"PE=mgH=mg\cdot L\text{ sin}\theta."

At the bottom this energy will turn into kinetic energy:


"KE=\frac{1}{2}mv^2."

Since


"PE=KE,\\
\space\\
mg\cdot L\text{ sin}\theta=\frac{1}{2}mv^2,\\
\space\\
v=\sqrt{2gL\text{ sin}\theta}=\\=\sqrt{2\cdot9.81\cdot916.6\cdot\text{sin}28.3^\circ}=92.3\text{ m/s}."

As we see, the final velocity does not depend on the mass the skier.